Two angles are complementary. The measure of the larger is six less than three times the measure of the smaller. Find the measures of the two angles.
Answer by checkley71(8403)

X+(3X-6)=90 OR X+3X-6=90 OR 4X=96 OR X=24 & 3*24-6=72-6=66 THUS 24+66=90